OpenStudy (anonymous):

Find the value of the variable(s). If your answer is an integer, leave it in simplest radical form. The triangle is not drawn to scale.

OpenStudy (anonymous):

\[a) x=17, y=34\sqrt{3}\] \[b) x=34, y=17\sqrt{3}\] \[c) x=34\sqrt{3}, y=17\] \[d) x=17\sqrt{3}, y=34\]

OpenStudy (anonymous):

I think its B.

OpenStudy (nikato):

No. I think you got the values of x and y mixed up. Also if you look at your answer, it doesn't make sense. Why would x, the leg, be longer than y, the hypotenuse